# RoTipBot: Robotic Handling of Thin and Flexible Objects using Rotatable Tactile Sensors

### v1

@@ Appendix A Configuration of FEM Simulation
Step: To replicate the actual dynamic contact process between the fingers and the paper, we chose the Dynamic Explicit Analysis method  [62] and divided the interaction into two steps. In the first step, the finger presses the paper; in the second step, the finger grasps the paper through rotation, with each step lasting one second. The output measured is the total force due to the frictional stress between the fingers and the paper, allowing us to compare the magnitude of the frictional forces required at different contact locations.
Interaction: We have defined two types of contacts: the contact between the fingers and the paper; the contact between the paper and the tabletop. For both types, normal contacts and tangential contacts are defined. Tangential contact is defined using a penalty friction formulation, with the friction coefficients set as 0.5 and 0.1, respectively [63].

### v2

@@ Appendix C Configuration of FEM Simulation
Step: To replicate the actual dynamic contact process between the fingers and the paper, we chose the Dynamic Explicit Analysis method and divided the interaction into two steps. In the first step, the finger presses the paper; in the second step, the finger grasps the paper through rotation, with each step lasting one second. The output measured is the total force due to the frictional stress between the fingers and the paper, allowing us to compare the magnitude of the frictional forces required at different contact locations.
Interaction: We have defined two types of contacts: the contact between the fingers and the paper; and the contact between the paper and the tabletop. For both types, normal contacts and tangential contacts are defined. Tangential contact is defined using a penalty friction formulation, with the friction coefficients set as 0.5 and 0.1, respectively [63].
@@
Appendix D RoTip Calibration
To achieve accurate tactile perception with RoTip sensors, two key components need to be calibrated: the camera intrinsic matrix, and the extrinsic offset $(o_{x},o_{y},o_{z})$ as shown in Fig. 5. Similar to the method in [22], we use a 7x8 checkerboard mounted on a board to get the sensor intrinsic with ROS calibration tools. During camera calibration, multiple sensor images are collected with different checkerboard poses to compute the camera intrinsic matrix $K$. Next, we calibrate extrinsic offsets $o_{x}$ and $o_{y}$ using tactile images from vertical contact with a flat surface, ensuring the contact area is centred on the fingertip. This step is treated as an optimisation problem aiming to minimise the distance between the predicted contact region’s centre and the fingertip’s top. Finally, we calibrate the offset $o_{z}$ along the camera’s $z$-axis by minimising the error between the predicted contact plane and the actual contact plane, using the tactile images collected at tilted angles, specifically 45 and -45. We chose not to adopt an end-to-end approach for contact surface prediction because it would require training a model for each individual RoTip sensor. Instead, by calibrating only the offset parameters $(o_{x},o_{y},o_{z})$ for each sensor, our approach simplifies the calibration process and enhances the model’s generalisation across RoTip sensors with varying installation errors.